Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (Kaz Borkowski at Wed Feb 13 18:38:42 MET 2002) Minor failures
No known problems other than those reported for this wavelength.

Medicina (Fiocchi Franco at Wed Feb 13 18:59:48 MET 2002) Minor failures
A lot of fog at start session. Some frequency settings of the procedure out of I.F. range for our VC01 & VC02.

\(Other\) (John Conway at Fri Feb 15 09:25:12 MET 2002) Minor failures
We observed till 03:23 then had a serious computer fault with the FS, with the machine hanging. After working on the computer in the morning we could continue observing from 10:06 to 16:40

Onsala ( at Fri Feb 15 09:25:37 MET 2002) Minor failures
Above entry for Onsala of course

Effelsberg (A. Kraus at Fri Feb 15 11:54:03 MET 2002) Minor failures
High parity errors.

Jodrell_Bank (Alastair Gunn at Mon Feb 18 11:04:01 MET 2002) Minor failures
Several antenna drive motor dropouts noted in log. No other known problems.

Westerbork (Tony Foley at Tue Feb 19 14:10:10 MET 2002) Minor failures
Scans after 16:25UT out of HA range. Beware possible 180deg phase jumps about 10:16, 14:34 and 14:50 UT